If O be the origin and the coordinates of P be (1, 2, -3), then find the equation of the plane passing through P and perpendicular to OP.

Grade:

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
The equation of the plane passing through P(1,2,-3) and perpendicular to OP
is r.n=a.n where r=xi+yj+zk
so x+2y-3z=1+4+9
x+2y-3z=14
Note: the above notations of r,n,a re vectors